OpenStudy (anonymous):

simplify the expression sqrt of 1/(3+8i)-(2-5i)

OpenStudy (anonymous):

\[\frac{ \sqrt{-1} }{ (3+8i)-(2+5i) }\]

OpenStudy (anonymous):

\[A. \frac{ 3+i }{ 10 } \] \[B. \frac{ -3+i }{ 10 }\] \[C. \frac{ 3-i }{ 10 }\] \[D. \frac{ (-3-i) }{ 10 }\]

OpenStudy (anonymous):

@Michele_Laino hey, are you busy at all m8?

OpenStudy (michele_laino):

first step, we have to simplify the denominator: \[\large \left( {3 + 8i} \right) - \left( {2 + 5i} \right) = 3 + 8i - 2 - 5i = \left( {3 - 2} \right) + \left( {8i - 5i} \right) = ...?\] please continue

OpenStudy (anonymous):

ok gimme a sec.

OpenStudy (anonymous):

1 + 3i

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

ok so how do we find the answer

OpenStudy (michele_laino):

so, we can write this: \[\Large \frac{{\sqrt { - 1} }}{{(3 + 8i) - (2 + 5i)}} = \frac{{\sqrt { - 1} }}{{1 + 3i}}\] now, we have to keep in mind that \(\sqrt{-1}=i\), so we have: \[\Large \frac{{\sqrt { - 1} }}{{(3 + 8i) - (2 + 5i)}} = \frac{{\sqrt { - 1} }}{{1 + 3i}} = \frac{i}{{1 + 3i}}\] next, second step: I multiply both numerator and denominator by \(1+3i\), so, I get: \[\Large \begin{gathered} \frac{{\sqrt { - 1} }}{{(3 + 8i) - (2 + 5i)}} = \frac{{\sqrt { - 1} }}{{1 + 3i}} = \frac{i}{{1 + 3i}} = \hfill \\ \hfill \\ = \frac{{i\left( {1 - 3i} \right)}}{{\left( {1 + 3i} \right)\left( {1 - 3i} \right)}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, we have to simplify these quantities: \[{i\left( {1 - 3i} \right)}=...?\] and: \[{\left( {1 + 3i} \right)\left( {1 - 3i} \right)}\]

OpenStudy (michele_laino):

oops.. I meant I multiply both numerator and denominator by \(1-3i\)

OpenStudy (anonymous):

ok so the first one becomes \[1i - 3i^2\]

OpenStudy (michele_laino):

yes! and \(i^2=-1\) so we get: \(i+3\)

OpenStudy (michele_laino):

then numerator is \(i+3\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh wait lol

OpenStudy (anonymous):

that's the answer

OpenStudy (anonymous):

hahahah thanks a lot

OpenStudy (michele_laino):

no, please wait: if I apply the distributive property, I get: \[\left( {1 + 3i} \right)\left( {1 - 3i} \right) = 1\left( {1 - 3i} \right) + 3i\left( {1 - 3i} \right) = ...?\] please continue

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok so then you get \[1-9i?\]

OpenStudy (michele_laino):

I got \(1-9i^2\)

OpenStudy (anonymous):

right my bad

OpenStudy (michele_laino):

which can be simplified to \(1-9(-1)=1+9=10\), since \(i^2=-1\)

OpenStudy (michele_laino):

so, the denominator is \(10\)

OpenStudy (anonymous):

ok so the answer is \[(3+i)/10\]

OpenStudy (michele_laino):

yes! :)

OpenStudy (anonymous):

thank you very much, have a nice day xx

OpenStudy (michele_laino):

:)